# AutoOcc: Automatic Open-Ended Semantic Occupancy Annotation via Vision-Language Guided Gaussian Splatting

### v2

@@ Abstract
Obtaining high-quality 3D semantic occupancy from raw sensor data remains an essential yet challenging task, often requiring extensive manual labeling. In this work, we propose AutoOcc, an vision-centric automated pipeline for open-ended semantic occupancy annotation that integrates differentiable Gaussian splatting guided by vision-language models. We formulate the open-ended semantic occupancy reconstruction task to automatically generate scene occupancy by combining attention maps from vision-language models and foundation vision models. We devise semantic-aware Gaussians as intermediate geometric descriptors and propose a cumulative Gaussian-to-voxel splatting algorithm that enables effective and efficient occupancy annotation. Our framework outperforms existing automated occupancy annotation methods without human labels. AutoOcc also enables open-ended semantic occupancy auto-labeling, achieving robust performance in both static and dynamically complex scenarios. All the source codes and trained models will be released.
@@ 1 Introduction
  • We present AutoOcc, a vision-centric automatic annotation pipeline that supports open-ended semantic occupancy label generation, based on vision-language guided differentiable reconstruction.
  • We devise VL-GS, an efficient and comprehensive scene representation for occupancy annotation. VL-GS integrates vision-language attention with visual foundation models, effectively handles dynamic objects over time, and enhances both spatiotemporal consistency and geometric detail.
  • AutoOcc gains notable improvements over the existing automatic occupancy annotation pipelines, even without relying on manual priors or LiDAR. Our method also demonstrates strong generalization and open-ended understanding capabilities.
@@ 2 Related Work
Scene Representation and Reconstruction. Efficient scene representation is the core to occupancy annotation. Dense voxel-based methods [24, 8, 25, 5] assign each voxel a feature vector, inevitably suffering from high computational cost due to redundant grids. As a compressed representation, BEV [16, 34, 6, 26] encodes 3D information on the ground plane, but struggles to capture diverse 3D geometry using flattened vectors. By implicitly modeling 3D space,  [63, 18, 15, 36] create a NeRF-style 3D volume to estimate scene occupancy. However, the continuous implicit neural fields struggle with modeling complex dynamic scenes, and dense sampling leads to redundant, memory-intensive operations. Most recently, 3D Gaussian splatting (3DGS) [20, 61, 38] has demonstrated its powerful capability in reconstruction, even for driving scenes [71, 48, 13]. By treating each vertex as a Gaussian,  [14] adopts a self-supervised approach for occupancy estimation but results in a dramatic increase in computational cost.
@@ 3.1 Vision-Language Guidance
We employ semantic attention maps to integrate and guide the acquisition of desired prior knowledge from vision-language models at the semantic level. Given a multi-view image sequence, we prompt the VLM [9] to consistently generate all possible object categories within each image. Specifically, we use the attention map generation method [1, 30] to compute and aggregate the attentions from transformer decoder, with $N$ output tokens $S=\{s_{1},\cdots,s_{N}\}$ and the attention tensor $A\in\mathbb{R}^{H\times L\times N\times N}$, with $H$ attention heads, $L$ layers:
@@
Semantic attention maps unveil category-related visual cues, which we subsequently leverage to guide the generation of semantic-aligned masks and depth information. Concretely, we input semantic attention maps as prompt cues into the the off-the-shelf segmentation models [69, 62], which then generates multiple masks within the region of interest. These masks are merged into instance-level candidate masks to fully delineate the targeted semantic regions. The mask with the highest similarity score to the embeddings of the semantic attention query is then selected.
In parallel, we employ semantic attention maps to guide depth estimation [37, 59] at the semantic level, decoupling background and foreground objects while excluding sky regions to avoid interference from infinite distances. We then aggregate depth information from multi-view images using semantic attention cues, where pixels within each region of interest yield a set of pseudo 3D point clouds that represents an individual instance.
@@
3.2 VL-GS
Although vision-language guidance provides valuable priors, 3D occupancy annotation still encounters three major challenges: 1) Semantic conflicts across multi-views make naïve 2D-to-3D projection prone to misalignment and ambiguity; 2) Errors in depth estimation lead to geometric distortions in 3D space; 3) Dynamic objects disrupt both spatial and temporal consistency in semantics and geometry.
@@ 3.2 VL-GS
Semantic-aware Scalable Gaussian.
$G(x)=e^{-\frac{1}{2}(\kappa-o)^{\top}\Sigma^{-1}(\kappa-o)}.$ (4)
The scaling factor is then adaptively adjusted based on the gradients of Gaussian values and the occupied range of the Gaussians. Notably, Gaussians of the same semantic category share similar scaling factor ranges, as objects with the same semantics exhibit comparable scales and geometries. As shown in Figure 2, semantic-aware scalable gaussians enable the representation of large background areas (e.g., buildings) with sparse gaussians at a larger scale, while capturing finer geometries (e.g., cyclist) with denser gaussians at a smaller scale.
Self-estimated Flow for Dynamic Objects.
Specifically, we model the translation of each Gaussian kernel $p$ from time $t$ to time $t+\Delta t$ as a flow vector $f$. Our goal is to minimize the point distances between object’s source points $U_{1}$ and target points $U_{2}$ to estimate the flow by applying Chamfer distance (CD) [12]. Since the same dynamic object is often represented by spatially adjacent Gaussians with the same semantics, we search for correspondences between paired points among the nearest Gaussian neighbors that share the same semantic:
where $p$ and $p^{\prime}$ are the position of Gaussian kernels with the same semantic at time $t$ and $t+\Delta t$, respectively. we define a dynamic indicator function between paired Gaussians to determine whether an object is in motion:
where $D$ is the average distance between paired Gaussians with the same semantic, $\rho$ is the dynamic threshold, $m$ denotes the number of Gaussian ellipsoids. The centroid position at the $i$-th frame is denoted by $o_{i}$. Subsequently, we aggregate all temporally paired Gaussians based on semantic attention map and motion cues.
Figure 3: Qualitative results of semantic occupancy annotation on Occ3D-nuScenes [49]. Our method achieves annotation accuracy and completeness comparable to human labeling, outperforming current multi-stage offline and self-supervised semantic occupancy ground truth generation pipelines. AutoOcc demonstrates good performance in capturing fine-grained geometry, ensuring semantic consistency, and handling temporal dynamics.
Geometric constraints from LiDAR.
Similar to [68, 67], a point $p_{i,t}$ in the LiDAR sweep $L_{t}$ is projected onto the frame $I_{t}$, and its initial semantic label can be obtained by $K^{-1}[R^{\top}\phi_{x,y,t}+T]$, where $(K,R,T)$ are the corresponding camera parameters and homogenous transformation matrix, and $\phi$ is the pixel-level semantic label. We aggregate the multi-frame of LiDAR points over time and compute the anchor centers $p_{c}=(x_{c}^{i},y_{c}^{i},z_{c}^{i})$. We then implement a geometry-aware loss to enforce the alignment of Gaussian ellipsoid distributions with the geometric priors of their corresponding semantic regions:
$L_{geo}=-\sum_{c=1}^{C}\sum_{i=1}^{M}\frac{1}{\|o_{c}(i)-p_{c}(i)\|_{2}^{2}},$ (7)
Cumulative GS-Voxel Splatting.
@@ 4 Experiments
Table 3: Zero-shot cross dataset performance on SemanticKITTI [2]. Other-veh. and moto-cyc. are short for Occ3D-nuScenes, other-vehicle, and motorcycles, respectively. Novel class refers to unseen semantics, while base class includes those seen during training. Metric mIoU-base denotes the mIoU computed solely on base classes.
@@ 4.1 Implementation Details
We use two benchmarks for evaluation: Occ3D-nuScenes, which is used to compare the performance of our method with other occupancy annotation methods for specific categories, while SemanticKITTI is used to assess the zero-shot capability across datasets and unseen categories. We set the resolutions of images as 900 × 1600 for Occ3D-nuScenes and 370 × 1226 for SemanticKITTI. During optimization, we scale the image size to 225 × 400 and double it every 300 steps until reaching the original resolution. Same as [30], we chose CogVLM-17B [52] with EVA2-CLIP-E [46] and Vicuna-7B-v1.5 [11] as the vision-language model. We follow GenerateU [29] to adopt CLIP [39] text encoder and map the generated categories to predefined categories in datasets for evaluation. We use the AdamW optimizer for optimization with an initial learning rate of 0.005. The learning rate for the position parameters decays every 250 steps with a decay rate of 0.98.
4.2 Performance Evluation and Analysis
@@
Point-based voxelization annotation pipelines directly use LiDAR with 3D annotations (semantic points and 3D bounding boxes) as input. SurroundOcc [55] performs mesh reconstruction and nearest neighbors algorithm to densify semantic points. OpenOcc [53] proposes the AAP pipeline to densify the voxel, followed by human post-processing to purify artifacts. For fair comparisons, we replicate these methods by replacing the manually annotated results with the semantic point clouds projected from VLMs. As shown in Table 2, our vision-centric method outperforms these pipelines that utilize LiDAR point clouds.
@@
Projecting annotated or generated 2D labels back onto 3D representation is a natural idea, which is further refined by several methods [58, 33]. However, these methods rely on pre-built 3D representations (e.g., point clouds or mesh) and employ multi-stage post-processing, including voting, filtering, and merging, to eliminate overlapping information. Undoubtedly, this strategy leads to the loss of crucial details and misalignment between semantics and representations. AutoOcc performs well against SAMPro3D [58] and OVIR-3D [33], both of which project the outputs of SAM [43] onto 3D point clouds. We also design a baseline that directly projects the results of VLM and SAM onto LiDAR point clouds (VLM-LiDAR) and voxelizes them into semantic occupancy. Table 2 shows that still demonstrates better performance, based on the deep integration of VLM guidance and differentiable reconstruction.
@@
Figure 3 and 4 shows that our method excels in semantic occupancy annotation, showcasing superior scene completeness, consistency, and dynamic object handling, even without the use of LiDAR. In extreme weather conditions (e.g., rain and nighttime), our method maintains robust performance, achieving annotation results comparable to or even surpassing manually labeled ground truth. For instance, in areas where ground truth is missing due to rain, AutoOcc successfully reconstructs both the geometry and semantics of the road surface.
@@ 4.4 Annotation Efficiency
Table 4 presents evaluations on representation characteristics and model efficiency. Notably, AutoOcc demonstrates an advantage in computational cost, delivering better performance with reduced memory requirements. In contrast, scene representations based on dense voxels and Point Cloud incur redundant computational costs. In addition, AutoOcc strikes a balance between efficiency and flexibility, enabling open-ended scene-aware occupancy reconstruction, supporting open-vocabulary semantic occupancy annotation, and requiring no human-labeled annotations.
Figure 5: Semantic occupancy of dynamics. AutoOcc accurately annotate semantic occupancy of dynamic objects, maintains spatiotemporal consistency, and infers occluded parts.
Table 5: Effect of each module in our method. SFM is short for the self-estimated flow module and SSG denots the employment of the semantic-aware scalable gaussians.
@@
4.5 Ablation Studies
We analyze the effect of self-estimated flow module for dynamic objects by disabling the clustering of dynamic objects and optimizing them together with static foregrounds. Figure 5 shows that the self-estimated flow module effectively mitigates the challenges of dynamic trailing and spatial occlusion in the annotation of occupancy. We further ablate the effect of LiDAR geometric priors and semantic-aware scalable Gaussians by either removing the $L_{geo}$ loss or replacing our SSG with vanilla Gaussians in our framework. The degraded results highlight the importance of these modules in constraining the shape and distribution of Gaussians, thereby enabling a more accurate reconstruction of the overall scene structure. More quantitative and qualitative results are available in the supplementary material.
@@ 5 Conclusion
In this paper, we propose AutoOcc, an vision-centric automated pipeline for open-ended semantic occupancy annotation that integrates differentiable Gaussian splatting guided by vision-language models. To facilitate scene understanding, we leverage VLMs and build an efficient and comprehensive scene representation for occupancy annotation. AutoOcc integrates vision-language attention with visual foundation models, effectively handles dynamic objects over time, and enhances both spatiotemporal consistency and geometric detail. Our framework achieves state-of-the-art performance on open-ended semantic occupancy annotation and performs favorably against other automated annotation pipeline, without using any human annotations.

### v3

@@ Abstract
Obtaining high-quality 3D semantic occupancy from raw sensor data remains an essential yet challenging task, often requiring extensive manual labeling. In this work, we propose AutoOcc, a vision-centric automated pipeline for open-ended semantic occupancy annotation that integrates differentiable Gaussian splatting guided by vision-language models. We formulate the open-ended semantic 3D occupancy reconstruction task to automatically generate scene occupancy by combining attention maps from vision-language models and foundation vision models. We devise semantic-aware Gaussians as intermediate geometric descriptors and propose a cumulative Gaussian-to-voxel splatting algorithm that enables effective and efficient occupancy annotation. Our framework outperforms existing automated occupancy annotation methods without human labels. AutoOcc also enables open-ended semantic occupancy auto-labeling, achieving robust performance in both static and dynamically complex scenarios.
@@ 1 Introduction
  • We present AutoOcc, a vision-centric automatic annotation pipeline that supports open-ended semantic 3D occupancy label generation, based on vision-language guided differentiable reconstruction.
  • We devise VL-GS, an efficient and comprehensive scene representation for semantic occupancy annotation. By integrating vision-language attention with visual foundation models, VL-GS effectively handles dynamic objects over time while enhancing both spatiotemporal consistency and 3D geometric detail in semantic occupancy.
  • AutoOcc gains notable improvements over the existing automatic occupancy annotation pipelines, even without relying on manual priors. Our method also demonstrates strong generalization and open-ended understanding capabilities across unseen categories and diverse scenes.
@@ 2 Related Work
Scene Representation and Reconstruction. Efficient scene representation is the core of occupancy annotation. Dense voxel-based methods [23, 8, 24, 5] assign each voxel a feature vector, inevitably suffering from high computational cost due to redundant grids. As a compressed representation, BEV [15, 33, 6, 25] encodes 3D information on the ground plane, but struggles to capture diverse 3D geometry using flattened vectors. By implicitly modeling 3D space,  [59, 17, 14, 35] create a NeRF-style 3D volume to estimate scene occupancy. However, the continuous implicit neural fields struggle with modeling complex dynamic scenes, and dense sampling leads to redundant, memory-intensive operations. Most recently, 3D Gaussian splatting (3DGS) [19, 57, 37] has demonstrated its powerful capability in reconstruction, even for driving scenes [67, 45, 12]. By treating each vertex as a Gaussian,  [13] adopts a self-supervised approach for occupancy estimation but results in a dramatic increase in computational cost.
@@ 3.1 Vision-Language Guidance
We employ semantic attention maps to integrate and guide the acquisition of desired prior knowledge from vision-language models at the semantic level. Given a multi-view image sequence, we prompt the VLM [9] to consistently generate all possible object categories within each image. Specifically, we use the attention map generation method [1, 29] to compute and aggregate the attentions from transformer decoder, with $N$ output tokens $S=s_{1},\cdots,s_{N}$ and the attention tensor $A\in\mathbb{R}^{H\times L\times N\times N}$, with $H$ attention heads, $L$ layers:
@@
Semantic attention maps unveil category-related visual cues, which we subsequently leverage to guide the generation of semantic-aligned masks and depth information. Concretely, we input semantic attention maps as prompt cues into the off-the-shelf segmentation models [65, 58], which then generate multiple masks within the region of interest. These masks are merged into instance-level candidate masks to fully delineate the targeted semantic regions. The mask with the highest similarity score to the embeddings of the semantic attention query is then selected.
In parallel, we employ semantic attention maps to guide depth estimation [36, 55] at the semantic level, decoupling background and foreground objects while excluding sky regions to avoid interference from infinite distances. We then aggregate depth information from multi-view images using semantic attention cues, where pixels within each area of interest yield a set of pseudo 3D point clouds that represent an individual instance.
3.2 VL-GS
Although vision-language guidance provides valuable priors, 3D occupancy annotation still encounters three major challenges: 1) Semantic conflicts across multiple views make naïve 2D-to-3D projection prone to misalignment and ambiguity; 2) Errors in geometry and depth estimation can lead to distortions and misplacements in 3D space; 3) Dynamic objects disrupt both spatial and temporal consistency in semantics and geometry.
@@ 3.2 VL-GS
Semantic-aware Scalable Gaussian.
$G(x)=e^{-\frac{1}{2}(\kappa-o)^{\top}\Sigma^{-1}(\kappa-o)}.$ (4)
The scaling factor is then adaptively adjusted based on the gradients of Gaussian values and the occupied range of the Gaussians. Notably, Gaussians of the same semantic category share similar scaling factor ranges, as objects with the same semantics exhibit comparable scales and geometries. As shown in Figure 2, semantic-aware scalable gaussians enable the representation of large background areas (e.g., buildings) with sparse gaussians at a larger scale, while capturing finer geometries (e.g., cyclists) with denser gaussians at a smaller scale.
Self-estimated Flow for Dynamic Objects.
Specifically, we model the translation of each Gaussian kernel $p$ from time $t$ to time $t+\Delta t$ as a flow vector $f$. Our goal is to minimize the point distances between the object’s source points $U_{1}$ and target points $U_{2}$ to estimate the flow by applying Chamfer distance (CD) [11]. Since the same dynamic object is often represented by spatially adjacent Gaussians with the same semantics, we search for correspondences between paired points among the nearest Gaussian neighbors that share the same semantic:
where $p$ and $p^{\prime}$ are the positions of Gaussian kernels with the same semantic at time $t$ and $t+\Delta t$, respectively. We define a dynamic indicator function between paired Gaussians to determine whether an object is in motion:
where $D$ is the average distance between paired Gaussians with the same semantic, $\rho$ is the dynamic threshold, and $m$ denotes the number of Gaussian ellipsoids. The centroid position at the $i$-th frame is denoted by $o_{i}$. Subsequently, we aggregate all temporally paired Gaussians based on the semantic attention map and motion cues.
Geometric constraints from LiDAR.
[…]
$L_{geo}=-\sum_{c=1}^{C}\sum_{i=1}^{M}\frac{1}{\|o_{c}(i)-p_{c}(i)\|_{2}^{2}},$ (7)
Cumulative GS-Voxel Splatting.
@@ 4 Experiments
Table 3: Zero-shot cross-dataset performance on SemanticKITTI [2]. Novel class refers to entirely new, unseen semantics in nuScenes, while base class includes those seen during training. All compared methods are trained on Occ3D-nuScenes and evaluated on SemanticKITTI. Metric mIoU-base denotes the mIoU computed solely on base classes from Occ3D-nuScenes.
@@
Figure 3: Qualitative results of semantic occupancy annotation on Occ3D-nuScenes [46]. Our method enables high-quality annotation of semantic 3D occupancy, capturing fine-grained geometry, structurally challenging regions, and dynamic objects across complex scenes.
@@ 4.1 Implementation Details
We use two benchmarks for evaluation: Occ3D-nuScenes, which is used to compare the performance of our method with other occupancy annotation methods for specific categories, while SemanticKITTI is used to assess the zero-shot capability across datasets and unseen categories. We set the resolutions of images as 900 × 1600 for Occ3D-nuScenes and 370 × 1226 for SemanticKITTI. During optimization, we scale the image size to 225 × 400 and double it every 300 steps until reaching the original resolution. We use the AdamW optimizer for optimization with an initial learning rate of 0.005. The learning rate for the position parameters decays every 250 steps with a decay rate of 0.98. Similar to [59, 66], we evaluate without the “other” and “other flat” classes. All experiments are carried out on 8 A100 GPUs.
4.2 Performance Evaluation and Analysis
@@
Point-based voxelization annotation pipelines directly use LiDAR with 3D annotations (semantic points and 3D bounding boxes) as input. SurroundOcc [51] performs mesh reconstruction and a nearest neighbors algorithm to densify semantic points. OpenOcc [49] proposes the AAP pipeline to densify the voxel, followed by human post-processing to purify artifacts. For fair comparisons, we replicate these methods by replacing the manually annotated results with the semantic point clouds projected from VLMs. As shown in Table 2, our vision-centric method outperforms these pipelines that utilize LiDAR point clouds.
@@
Projecting annotated or generated 2D labels back onto a 3D representation is a natural idea, which is further refined by several methods [54, 32]. However, these methods rely on pre-built 3D representations (e.g., point clouds or meshes) and employ multi-stage post-processing, including voting, filtering, and merging, to eliminate overlapping information. Undoubtedly, this strategy leads to the loss of crucial details and misalignment between semantics and representations. AutoOcc performs well against SAMPro3D [54] and OVIR-3D [32], both of which project the outputs of SAM [41] onto 3D point clouds. We also design a baseline that directly projects the results of VLM and SAM onto LiDAR point clouds (VLM-LiDAR) and voxelizes them into semantic occupancy. Table 2 shows that AutoOcc still demonstrates better performance, based on the deep integration of VLM guidance and differentiable reconstruction.
@@
Figure 3 and 4 show that our method excels in semantic occupancy annotation, showcasing superior scene completeness, consistency, and dynamic object handling, even without the use of LiDAR. In extreme weather conditions (e.g., rain and nighttime), our method maintains robust performance, achieving annotation results comparable to or even surpassing manually labeled ground truth. For instance, in areas where ground truth is missing due to rain, AutoOcc successfully reconstructs both the geometry and semantics of the road surface.
@@ 4.4 Annotation Efficiency
Table 4 presents evaluations on representation characteristics and model efficiency. Notably, AutoOcc demonstrates an advantage in computational cost, delivering better performance with reduced memory requirements. In contrast, scene representations based on dense voxels and Point Clouds incur redundant computational costs. In addition, AutoOcc strikes a balance between efficiency and flexibility, enabling open-ended scene-aware occupancy reconstruction, supporting open-vocabulary semantic occupancy annotation, and requiring no human-labeled annotations.
Figure 5: Semantic occupancy of dynamics. AutoOcc accurately annotates the semantic occupancy of dynamic objects, maintains spatiotemporal consistency, and infers occluded parts.
4.5 Ablation Studies
We analyze the effect of the self-estimated flow module for dynamic objects by disabling the clustering of dynamic objects and optimizing them together with static foregrounds. Table 5 and Figure 5 show that the self-estimated flow module effectively mitigates the challenges of dynamic trailing and spatial occlusion in the annotation of occupancy. We further ablate the effect of semantic-aware scalable Gaussians and LiDAR geometric priors by replacing the scalable Gaussians and removing the $L_{geo}$ loss from our framework. The degraded results highlight the importance of geometric priors in constraining the shape and distribution of Gaussians, leading to more accurate occupancy reconstruction.
Table 5: The effect of each module in our method. SFM is short for the self-estimated flow module, and SSG denotes the employment of the semantic-aware scalable Gaussians.
@@ 5 Conclusion
In this paper, we propose AutoOcc, an vision-centric automated pipeline for open-ended semantic 3D occupancy annotation that integrates differentiable Gaussian splatting guided by vision-language models. To facilitate scene understanding, we leverage VLMs and build an efficient and comprehensive scene representation for occupancy annotation. AutoOcc integrates vision-language attention with visual foundation models, effectively handles dynamic objects over time, and enhances both spatiotemporal consistency and geometric detail. Our framework achieves state-of-the-art performance on open-ended semantic occupancy annotation and performs favorably against other automated annotation pipeline, without using any human annotations.
Acknowledgment
This work was supported by National Key R&D Program of China (Grant No. 2022ZD0160305). This work was also a research achievement of Key Laboratory of Science, Technology, and Standard in Press Industry (Key Laboratory of Intelligent Press Media Technology). Ming-Hsuan Yang was supported in part by the Institute of Information & Communications Technology Planning & Evaluation (IITP) grant funded by the Korean Government (MSIT) (No. RS-2024-00457882, National AI Research Lab Project).